40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 1 of 349  Not logged in ELOG logo
ID Date Author Type Category Subject
  17611   Thu Jun 1 10:03:47 2023 YehonathanUpdateSUSETMX Coil driver upgrade

Thanks. This makes the plan clear.

- Please follow the attached rack plan for 1X9

- The immediate aim is the recovery of the suspension operation without acromag. Focus on it for now, and then move on to the Acromag modification.

-----

> Take out 2 new coil drivers and replace their output resistance with appropriate resistors (What should they be?)

We are supposed to have the coil drivers with the output Rs already replaced. But it is a great idea to check if they are really done or not.

> Test the new coil driver on the bench.

I hope this is not necessary, as the coil drivers / sat amps were already tested before...

> Change the powering of the Acromag chassis so that we can operate it with a single Sorensen.

Good idea. If we ended up having an unnecessary Sorensen, we can remove it and add it to the stock.

 

Attachment 1: rack_plan.pdf
rack_plan.pdf
  17610   Thu Jun 1 08:34:06 2023 PacoUpdateCDSITMX UL coil and BS oplev rms

ITMX / BS local damping issues  -- investigation in progress

[Paco, Yuta]

Background

We noticed an increase in PD variance in a few optics since last Friday (05/26, circa 23:00 UTC) and decided to investigate using dataviewer. There are a few observations made by Yuta:

  • BS -- OpLev rms increase, OSEMs are nominal
    • May 26 ~22:00 UTC, oplev PIT/YAW rms increased suddenly
    • May 28 ~ 17:00 UTC, optic tripped (model crash)
    • May 29 ~21:00 UTC, optic restored (model restart), oplev PIT/YAW rms still increased
  • ITMX -- OpLev rms increase, OSEMs are not nominal (specifically UL)
    • May 26 ~ 22:00 UTC, oplev glitch but rms stays nominal
    • May 28 ~17:00 UTC, optic tripped (model crash)
    • May 29 ~21:00 UTC, optic restored (model restart), oplev PIT/YAW rms increased suddenly

BS oplev alignment

We thought it would be worth checking the BS oplev table for any alignment issues, since playing with the loop gains (OL_PIT/YAW_GAIN) didn't seem to have an effect. We quickly found two issues:

  1. Clipping on lens before QPD: which was fixed by alignment and recentering of the lens aperture.
  2. Lens AR coating was dumping almost 50% of the power: which was fixed by switching from a f=62.9mm AR.33 coated (for 1064 nm) plano convex lens to an f~ 100 mm uncoated plano convex thorlabs singlet (LA1509).

These changes are always good to implement, but the increased BS oplev rms issue remains.

Coil output test

Under the premise that local damping is not working correctly, we decided to test coil outputs at the coil driver test monitor outputs and also before (at the dewhitening monitor outputs). Looks like BS_ULCOIL is not able to drive a large enough voltage compared with the other Coils. This is consistent with the very weak optic "kicks" we get on BS when using this coil.

 

After this issue, we restarted all models.

 

Attachment 1: Screenshot_2023-05-31_16-49-26_BS.png
Screenshot_2023-05-31_16-49-26_BS.png
Attachment 2: Screenshot_2023-05-31_16-51-44_BSoplev.png
Screenshot_2023-05-31_16-51-44_BSoplev.png
Attachment 3: Screenshot_2023-05-31_16-39-28_ITMX.png
Screenshot_2023-05-31_16-39-28_ITMX.png
Attachment 4: Screenshot_2023-05-31_16-56-47_ITMXoplev.png
Screenshot_2023-05-31_16-56-47_ITMXoplev.png
  17609   Wed May 31 16:26:28 2023 YehonathanUpdateSUSETMX Coil driver upgrade

{Mayank, Yehonathan}

In order for the ETMX watchdog to keep working while we take the Acromag chassie out we are going to replace the PD_MONs in the variance calculation channels with fast OSEM PD channels/

We made a backup to the c1auxex db file called ETMXauxbak.db in the c1auxex folder.

Now, we are struggling with adding channels to the db file. Once we figure this out we will try to read the fast SENS_IN1 channels.

In the meanwhile this is a broadview (and maybe naive) plan for the coil driver upgrade:

-1. Take out 2 new coil drivers and replace their output resistance with appropriate resistors (What should they be?)

0. Test the new coil driver on the bench.

1. Once a temporary fast-channel-based watchdog is setup and tested, Acromag chassie is taken out to the EE bench.

2. Replace the old DB37 feedthroughs with DB9s according to the ETMY plan. Check if there are extra channels that exist in c1auxex but not in c1auxey that need to be preserved.

3. Wire the acromag units to the DB9s. Add optical isolators and wire the EnableMon and Enable fast channels through them.

4. Change the powering of the Acromag chassie so that we can operate it with a single Sorensen.

5. Add 5 EnableMon channels. We will probably need anothe BIO Acromag as there is only 1 in current chassie according to the old wiring table.

5. Bench-test the Acromag chassie channels.

6. Add 5 new EnableMon channels to the db file.

7. Install Acromag chassie, coil drivers, SatAmps and HV drivers and the connect labeled db9 cables between them.

8. Test the new coil drivers online.

Anything else?

 

  17608   Wed May 31 12:08:07 2023 ranaUpdateBHDSensing matrix model

it is great to see a sensing matrix without 900 digits of precision!

for choosing what sensor to use, we don't necessarily care about W/m, but more like the equivalent noise of the sensor in m/rtHz, taking into account the real noise floor. In that case, we would possibly rotate the demod phase to maximize the SNR rather than maximize the S or minimize the N.

  17607   Wed May 31 10:23:40 2023 YehonathanUpdateBHDSensing matrix model

I calculated the sensing matrix for PRMI carrier using the Finesse model (git updated) using MAXTEM=2. PRG is calculated to be 11.14, consistent with observations.

LO Phase is chosen such that it maximizes MICH signal on BHD_DIFF.

The RFPDs were assumed to have a demodulation angle that maximizes the signal they are intended to sense (using MAXTEM=2).

That is,

BH44/55 maximized for HPC

REFL11 maximized for PRCL

AS55 maximized for MICH

Values are in uW/nm

  MICH PRCL HPC
BHD_DIFF
50.00
73.67
0.68
BHD_SUM
17.28
440.80
3.81e-14
BH44
0.38
4.22
3.1e-2
BH55
0.4
3.17
3.3e-2
REFL11_I 1.9e-2
13.46
0

AS55_Q

7.6e-2 4.0e-2 0

Some interesting numbers here. First, BHD_SUM is sensitive to MICH. It's not surprising because PRM reflects the MICH signal into POP.

Also interesting, BHD_SUM is super sensitivef to PRCL. Much more than REFL11. We can use it to enhance the PRCL lock.

Unfortunately, although BH44/55 are sensitive to HPC (LO phase), they are swamped by MICH and PRCL. This issue needs to be addressed in order to gain robust LO Phase locking in PRMI.

  17606   Mon May 29 11:04:13 2023 PacoUpdateCDSc1sus2 all FE models crashed spontaneously again

c1sus2 crashed again. Following 40m/17335, I fixed it by running

controls@c1sus2:~$ rtcds restart --all

"global diag reset" made all FE STATUS green burt restored at 2023/May/28/00:19 for c1sus2 models, and watchdogs reset for BHD optics and now all look fine.


Other optics, including MC1, MC2, and MC3 were not damped, so maybe c1sus crashed too I also ran

controls@c1sus:~$ rtcds restart --all

"global diag reset" made all FE STATUS green, burt restored at 2023/May/28/00:19 for c1sus models and watchdogs reset for all other suspensions.


TT1 and TT2 were not responding, and the DAC monitors were frozen... so I ran

./opt/rtcds/caltech/c1/Git/40m/scripts/cds/restartAllModels.sh


Wed May 31 14:58:06 2023 UPDATE: I forgot to log that ITMX, ITMY and BS oplevs were centered after some nominal alignment was recovered.

  17605   Fri May 26 15:04:12 2023 JCUpdatePSLPSL Fans Replaced

We Changed The Fans on the PSL

To start, the part Koji ordered is 259-1818-ND from DigiKey. This is a Maglev fan from SUNON that should give us less noise. We have 3 spare replacement fans in case these go bad which are stored in the ___ Cabinet along the Y arm (This will be updated once I find a suitable storage spot for the part.)

Starting the replacement process.

Removing the PSL 

    1. Our first step to doing this was to prepare for removing PSL. We began by doing a 60s MC WFS Relief. This will allows us to turn off WFS and close the PSL Shutter next. This is to prevent a large kick once we place the PSL back in its place. 
    2. Went to the PSL and mushed the off button and turned the key. After this, begin by removing the external fan which is shown in elog 17595. After, continue by unplugging the cables from the back beginning with the power cable. Attachment #1 shows the original positions of the cables connected before removing any. Keep in mind, DO NOT TO TOUCH THE KNOBS. If the inputs are changed, this will throw off the beatnotes of the AUX lasers.
        a. There is a black plug at the bottom with a screw hat is hard to reach. Be very patient taking this off because the position of the cable blocks a screqwdriver from untightening the screw. 
        b. A second person should be on the other side of the table to push the module back into arm's reach. Also to make sure the module does not slide back and fall. 
    3. After removing the module, bring into the control room and place onnto the workbench. Make sure all of the red lights are off and the PSL table is closed properly.

Changing the Fan 
    
    
    1. We removed the top cover of the module and opened it all up. Similar to what is shown in elog 17452.
        a. Keep in mind to wear a grounding wristband when working on this.
    2. After removing the old fans and attempting to install the new ones, the holes did not line up correctly (Shown in Attachment #2). To accomodate for this, we used 6-32 screws which gave us just enough slack to fit in all 4 corners.
    3. Ones the fans were bolted down onto the aluminum plate, I soldered the cables to connecting the fan cable to the cables those of the original PSL fans.
    4. Next I used heat shrinks to cover the bare soldered areas and placed the fans into the module. 
    5. We tested the fans by plugging intp the PSL and turning the key. The fans turned on nicely and we proceeded to put it module back together.

Placing the PSL back in its place. 

    1. We place the PSL back into its original spot and began to connect the cables. Make sure the Power cables is put in LAST. 
    2. After the module was put back, we DID NOT put the external fan back into its place. This is to see if the fans which were installed are good enough to maintain the PSL.
    3. Turn the key and press the on Button.

The noise from the external fan is no longer appearing as shown in attachment #3. The PSL has been on for ~2 hrs now and has not turned off. It seems that the fans are doing their jobs well. 

 

Attachment 1: IMG_5650.jpeg
IMG_5650.jpeg
Attachment 2: IMG_5653.jpeg
IMG_5653.jpeg
Attachment 3: Screenshot_2023-05-26_16-25-54.png
Screenshot_2023-05-26_16-25-54.png
  17604   Fri May 26 11:16:46 2023 AnchalSummaryPEMTemperature sensing circuit with AD590

I wanted to mention here that I have a printed circuit board design (LIGO-D1800304) for using AD590 as temperature sensors. I believe printed boards and all required components are stored on the wire shelf in the WB EE shop on a box labeled with this DCC number. This circuit is also meant to be read by Acromag, maybe it can come of some use to you. You can check out the use in CTN lab in WB near the south-east end of the table.

 

  17603   Thu May 25 14:40:03 2023 PacoUpdateCalibrationMeasured the PSL wavelength

The PSL wavelength is 1064.5068 +- 0.0015 nm

[Paco, Yehonathan]

With yehonathan's help, we borrowed a WS6-200 wavemeter and calibrated the PSL wavelength using the frequency doubled beam at the PSL table. For this we gathered a FC/APC to FC/PC fiber, cleaned both ends, and launched ~ 6 uW of power into the wavemeter (Attachment #1 shows the setup). The vacuum wavelength was measured to be 532.2534 nm, (see picture in Attachment #2) implying a PSL wavelength of 1064.5068 nm. This is not too far from my previous estimate! The wavemeter claims a "200 MHz accuracy" so we used this as a standard error to estimate the uncertainty of  2 * 0.7 pm @ 532 nm = 1.5 pm @ 1064 nm. This leaves us with 1.4 ppm of relative wavelength error in the ALS based calibration.

  • It should be noted that the Nd:YAG temperature was set to 30.615(5) degrees (?) and the injection current to 2.100 Amps for this measurement.
Attachment 1: PXL_20230525_205542591~2.jpg
PXL_20230525_205542591~2.jpg
Attachment 2: PXL_20230525_205332570.jpg
PXL_20230525_205332570.jpg
  17602   Thu May 25 13:38:28 2023 advaitSummaryPEMExisting temperature control hardware

Over the past few days I have been trying to understand the existing sensor and heater related hardware by manual inspection and combing elogs. From what I understand, I believe a lot of the hardware was built from scratch by Kira in 2017-18. She put in place the insulation, built the sensor and heater circuits and installed and interfaced everything with EPICS. This let her successfully do step response tests and also execute PID control of the temperature of the can. As suggested by Paco, I am trying to summarise my findings in this elog. This is the block diagram of the overall system.

In the lab I was able to locate two identical temperature sensor boards, one for monitoring ambient temps and another for the can temp, which has a sensor attached to the inner seismometer can. Attachment 1 shows the actual board.

The schematic for the temperature sensor can be found here. It was later clarified that she is using AD590 and not the AD592. I tested the ambient sensor board by hooking it up to an oscilloscope and heating the sensor up. It seems to work like she described, and the output voltage goes down with heat. I will later figure out interfacing it with an ADC and calibrating it if I cannot find Kira's old work related to this. The setup there also has +/-15V Sorensen's for power (I probed these and they work), a BNC for temp readout and a bunch of other wiring for the heater. This wiring comes from the nearby rack which should have the heater circuit, but I haven't yet been able to locate it because there is too much stuff.

Coming to the heater circuit - I found the schematic for it here. I am still unsure about what the power range it can operate over, but this plot seems to imply it saturates at around 55W. This circuit underwent many different iterations so I am unsure what load resistance was actually used finally and it is not clear from the elogs. It might be a pair of heaters combined in series or parallel. Additionally, this elog by Shruti from June 2018 implies that the original circuit that Kira made ended up breaking at several different points. She said there were attempts to fix it, but I cannot find any updates after that and I think her SURF ended. I don't know the current status of the heater circuit. I am not sure where it is stored.

I was hoping to utilise the old EPICS channels set up by Kira for heater control as well as sensor readout to verify if everything was still functional and initially just read out the temperature of the can and find out how much it fluctuates. I planned to later also try to replicate the PID results. But Paco explained how the Acromag system worked and told me that utilising the old channels will not be possible right now, as all the channels have been used up for more critical tasks and none of the old interfaces that Kira had set up can be used now, even if the hardware was fine. 

Also, the can temperature sensor board has a wobbly solder joint at one of the power connectors, and we cannot move it to the lab because the sensor is anchored to the inner can. There is no power indicator LED to signal if the connection is secure so its kind of unusable right now. Paco and JC told me that soldering near the X end is not possible because the fumes would harm the optics. We will look for solutions to this. One option would be to make an entirely new board and keep a mechanical connector for the sensor. 

  17601   Wed May 24 17:36:25 2023 PacoUpdateBHDBH44_I content and PRC alignment

BH44 is sensitive to PRC alignment noise

[Paco, Yuta]

We investigated the content of BH44 demodulated signals under PRMI configuration. We had a few ideas of what was being sensed by BH44_I but we wanted to test this. Attachment #1 shows a timeseries screenshot of the DCPDs and BH44 error signals during PRMI lock stretch. It is pretty clear how BH44_I is sensing the same as REFLDC. To understand what REFLDC is sensitive to, we locked PRY (this is like having a lossy PRC) and looked at REFLDC, and BH44 error signals again. When PRY is aligned nicely, BH44 error signals show clean LO fringes and we could lock LO_PHASE stably (Attachment #2). Dithering the PRM YAW at 0.5 Hz (amplitude of 150 counts) is sensed by the REFLDC output, so we can attribute its fluctuations to the PRC misalignment (Attachment #3). Now we saw that the zero crossing of the homodyne phase angle changes following REFLDC, and LO_PHASE could not be locked stably. These suggest that alignment of PRC is sensed by BH44, and we might need alignment control to stably lock LO_PHASE in PRMI.
To get the idea of what is causing alignment fluctuations of PRC, we checked the spectrum of SUSPIT/YAW of PRM, PR2, PR3, BS, ITMX, and ITMY. It was not clear what is causing REFLDC fluctuations. (But we found that ITMX and ITMY has huge bounce mode at 16.2 Hz; see Attachment #4).

Next:
 - Check FINESSE to see what BH44 sees. PRCL? PRG?
 - Commission REFL WFS for alignment control of PRC?
 - Commission dither loops (add option to demodulate PRCL, modulate PR2 and PR3) for alignment control of PRC?
 - Check RF spectrum of BH44 and RF LO for 44 MHz (sidebands other than 44 MHz might be contaminating the signal).
 - Check ITMY scattering. Dither ITMY in YAW and check BH44.
 - Move on to PRMI sideband BHD

Attachment 1: Screenshot_2023-05-24_17-38-22_PRMI_REFLDCvsBH44.png
Screenshot_2023-05-24_17-38-22_PRMI_REFLDCvsBH44.png
Attachment 2: Screenshot_2023-05-24_17-40-58_PRY_REFLDCvsBH44.png
Screenshot_2023-05-24_17-40-58_PRY_REFLDCvsBH44.png
Attachment 3: Screenshot_2023-05-24_17-44-49_PRY_REFLDCvsBH44_PRM0.5Hz.png
Screenshot_2023-05-24_17-44-49_PRY_REFLDCvsBH44_PRM0.5Hz.png
Attachment 4: Screenshot_2023-05-24_17-46-13_ITMSBOUNCE.png
Screenshot_2023-05-24_17-46-13_ITMSBOUNCE.png
  17600   Wed May 24 13:19:28 2023 PacoUpdateBHDBH44 and BH55 dc transimpedance modified

We lowered the BH44 and BH55 DC transimpedances to ~ 50 V/A

[Paco, Yuta]

Background

When locking the homodyne phase angle using BH44_Q or BH55_Q error signals, we notice the orthogonal quadrature (BH44_I, BH55_I) sometimes appears too noisy. The origin of this useless signal is not known, but we have recently attenuated these beams by placing ND filters before the two RFPDs to avoid saturation effects which become obvious when we lock PRMI. We decided to investigate further by the following tests:

  • Remove ND filters and lower the DC transimpedances to ~ 50 V/A
  • Check for scattering from suspended optics, e.g. by injecting a line at ITM PIT/YAW and look at the BH44/BH55 demodulated spectra
  • Check for PRCL sensing by BH44/BH55, e.g. by measuring the transfer function and/or running a simulation in finesse.
  • Check the RF spectra for the signals entering the IQ demod boards (including the 44 and 55 MHz LOs).

DC transimpedance modifications

The first thing we did was change the DC transimpedances of both RFPDs. After removing them from the table, we checked the schematics for 40m RFPDs on the wiki. The DC transimpedance for these gold RFPDs (D980454-v1-C) is estimated as (R22*(1+R13/R23), where these resistors are located around the follower and non-inverting amplifier stages along the DC output traces. After opening the two RFPDs and taking photos of the circuits before any changes (Attachment #1-2), we estimated the DC transimpedances from the measured values for R22, R23 and R13 and summarized them below:

Before R13 R22 R23 Est. DC transimpedance
BH44 8.2 kOhm 10.4 Ohm 99.9 Ohm ~ 864.05 V/A
BH55 99.9 kOhm 12.6 Ohm 102.7 Ohm ~ 12.26 kV/A

The changes were made on R13 (photos in Attachments #2-3) and the final values summarized below:

Before R13 R22 R23 Est. DC transimpedance
BH44 402.4 Ohm 10.4 Ohm 99.9 Ohm ~ 52.29 V/A
BH55 309.5 Ohm 12.6 Ohm 102.7 Ohm

~ 50.57 V/A

All changes have been summarized and recorded in the wiki. The ND filters were set to 0.04 (minimum attenuation) and RFPDs reinstalled.

Next steps:

Continue investigating these items:

  • Remove ND filters and lower the DC transimpedances to ~ 50 V/A
  • Check for scattering from suspended optics, e.g. by injecting a line at ITM PIT/YAW and look at the BH44/BH55 demodulated spectra
  • Check for PRCL sensing by BH44/BH55, e.g. by measuring the transfer function and/or running a simulation in finesse.
  • Check the RF spectra for the signals entering the IQ demod boards (including the 44 and 55 MHz LOs).
Attachment 1: PXL_20230524_191008053~2.jpg
PXL_20230524_191008053~2.jpg
Attachment 2: PXL_20230524_192525594~2.jpg
PXL_20230524_192525594~2.jpg
Attachment 3: PXL_20230524_193122794~2.jpg
PXL_20230524_193122794~2.jpg
Attachment 4: PXL_20230524_194225616~2.jpg
PXL_20230524_194225616~2.jpg
  17599   Wed May 24 11:50:34 2023 JCSummarySEISTACIS

[Mayank, JC]

STACIS Vibration Cancellation Isolators contain a PCB board called the Compensator board. These are boards that "allow you to control the feedback to the PZT stacks." The board is essentially held up by the 2 rear plugs and 3 screws on the front rim (Highlighted in Yellow). You don't need to disassemble the entire islotar to remove this, you just need to loosen the tension screws and pull out softly. Attachment #4 shows the seat made for the board to be placed, highlighted in Green. When placing the board back into its original position make sure in slips into this slit nicely. The extender/compensator board is part of the STACIS Isolator and each Isolator should have this. Each board seems to haver it's own serial # but not an actual part #. On this board, the S/N is 033509.

Attachment 1: IMG_5633.jpeg
IMG_5633.jpeg
Attachment 2: IMG_5634.jpeg
IMG_5634.jpeg
Attachment 3: IMG_5638.jpeg
IMG_5638.jpeg
Attachment 4: IMG_5635.jpeg
IMG_5635.jpeg
  17598   Fri May 19 15:25:00 2023 MayankUpdatePSLPSL tripped - removed internal fans

We removed the PSL controller internal (broken) fans after it tripped due to overheat.

Background

[Mayank, Radhika]

While aligning Xarm I noticed sudden loss of beam. On Radhikas suggestion we cheked the PSL and found out that the PSL controller was in off state (No lights on front and back panel). We restored the situation by unplugging and replugging the power cord. The PSL worked fine for a few minutes (~ 30 ) and then tripped again.This time the front panel  OFF light was on . See attached image (Attachment #1).

Fix

[Paco, Mayank, Koji-remote]

We disconnected the PSL controller and took this opportunity to investigate the controller's internal cooling mechanism. After disassembling the top panel of the chassis, we saw there are two SUNON - KD1205PHB2 fans meant to run at 12 VDC (1.7 W) connected to the bottom pcb inside the controller. After disconnecting them from this board, we tested them with an externally supplied dc voltage and confirmed they no longer worked. We noted the cooling mechanism is based on a long aluminum heat sink to which most ICs are attached, and the fans are meant to provide airflow towards the rear aperture on the chassis. We followed Koji's suggestion and for now, removed the damaged components (detailed pictures of this operation have been posted in a google photos album elsewhere) to allow heat to flow out more easily.  We reassembled the controller chassis and reinstalled it with the external fan providing the necessary airflow to prevent the unit from tripping again due to overheating. Then we turned on PSL and recovered PMC and IMC locks.

Aftermath

We took a C1:IOO-MC_F_DQ trace after this work to confirm our earlier findings; the trace is attached in Attachment #2. The noise bumps are present as expected. This is still not a desirable configuration so next step would be replacing the external fan, or even better, find the appropriate spare of the internal units and berid the external one.

Attachment 1: 20230519_123659.jpg
20230519_123659.jpg
Attachment 2: afterfans_PSLcontroller_Screenshot_2023-05-19_15-48-04.png
afterfans_PSLcontroller_Screenshot_2023-05-19_15-48-04.png
  17597   Fri May 19 13:25:03 2023 KojiUpdatePSLMCF Noise

This is super! And now is the time to replace the internal fan!

 

  17595   Fri May 19 09:23:58 2023 PacoUpdate MCF Noise

The fan behind the PSL controller is injecting excess band limited noise angry

Background

While doing noise hunting to improve the BHD lock stability, we noticed peculiar noise bumps in the BH44 error point near (but not exactly at) the even line-harmonics (for example 120 Hz, 240 Hz, ...). Other channels such as C1:HPC-BHDC_SUM, C1:LSC-PRCL_IN1, or even C1:LSC-XARM_IN1 didn't show these features, so we looked at C1:IOO-MC_F_DQ (which represents the free running laser noise above 100 Hz) and to our surprise found this excess noise!

Noise hunting around PSL

Since this noise is present upstream, we decided to hunt around using C1:IOO-MC_F_DQ. We set up a diaggui measurement to do some "live demodulation" as suggested by Koji in order to understand the nature of this noise. In order to get some "video bandwidth" we set up a power spectrum measurement from 114 Hz to 140 Hz (to monitor the usual 120 Hz line noise peak) with a bandwidth of 1 Hz. A single exponential average gave us the 1 Hz narrow spectrum in "real time", from which we noticed its nonstationary character. The band limited excess noise is the result of a peak hovering in the range of 125 to 131 Hz. With this diagnostic set up, we started hunting for its source.

  1. We checked the fan behind the PSL controller (Attachment #1). 
    1. After disconnecting the molex powering it with +15 VDC the noise dissappeared!

To show the impact in the complete noise spectrum, we took a 10 fixed average measurement with and without the fan being on. The result is in Attachement #2. The spectra are shown along with their rms, which is significantly reduced when the fan is off near the 100 Hz frequency band (where these bumps appear). Anyways, we have left the fan on because the PSL controller needs it so the problem remains, but we have at least identified the source.

Attachment 1: PXL_20230519_162346524.jpg
PXL_20230519_162346524.jpg
Attachment 2: MC_Fnoisehunting_Screenshot_2023-05-19_09-27-03.png
MC_Fnoisehunting_Screenshot_2023-05-19_09-27-03.png
  17594   Wed May 17 12:09:11 2023 YehonathanUpdateElectronicsPreping for new coil drivers commissioning in 1X4

It's time to start commissioning the new coil drivers. The Acromag box is already there but it needs to be modified.

The idea is to take the channel configuration from C1AUXEY that controls a single suspension - ETMY and apply it for each suspension in 1X4, that is PRM, BS, ITMX, ITMY.

For this, we need the following items

Item # per SUS # Total
DB9 Front panels 2 8
Optical isolators 4 16
DB9M feedthrough 6 24
DB9F feedthrough 5 20
DB9 Cables ? ?
EnableMON BIO channels 5 20

Currently, there are 8 spare BIO channels in the existing Acromag, we will need 12 more which requires a new BIO Acromag.

Wiring feedthrough spreadsheet coming soon.

 

 

  17593   Tue May 16 08:42:14 2023 PacoUpdateASCFPMI differential ASC - II

We aligned the AS WFS and measured some angular motion from ETMX


This work is done after IMC WFS offsets are offloaded and the WFS loops are turned off. Then, the flipper mirror M4 in the IMC WFS path is flipped up to feed the AS beam into its path. We also connected the AS WFS demod boards for both heads.

[JC, Paco]

We continued to get the AS WFS going. We added two lenses with the help of JC (f1~500 mm and f2~1000 mm) to have control over the minimum waist and its position downstream. In practice, we placed the lenses such that the beam waist lands close to the last lens before WFS1, giving a reasonable beamspot at both WFS1 WFS2 heads. These changes are reflected by the picture on Attachment #1. The beam sizes will sure need careful optimization to sense the Differential/Common ARM angular motion due to ITMs and ETMs, but for now we decided to move forward.

[Mayank, Paco]

We locked FPMI and use the WFS LOCKINs to inject a 21.173 Hz line into ETMX-PIT. Then, we look at the AS11-WFS signals using diaggui to identify any measurable signal. We see the line coming into a combination of channels C1:ASC-DHARD_PIT_IN1, C1:ASC-DHARD_YAW_IN1, C1:ASC-DSOFT_PIT_IN1, and C1:ASC-DSOFT_YAW_IN1 as shown by Attachment #2.

Attachment 1: PXL_20230515_173719320~2.jpg
PXL_20230515_173719320~2.jpg
Attachment 2: ASWFS_test0_Screenshot_2023-05-16_18-38-39.png
ASWFS_test0_Screenshot_2023-05-16_18-38-39.png
  17592   Mon May 15 18:34:05 2023 PacoSummaryLSCLocked PRMI in carrier and LO phase with BH44 and achieved handoff to BHDC_DIFF

[Yuta, Paco]

We locked PRMI-BHD (LO phase was controlled using BH44_Q) cool


Today we worked with PRMI carrier locked using AS55_Q and REFL11_I (to control MICH and PRCL respectively).

PRCL and MICH displacement calibration

We calibrated the PRCL and MICH displacements. This time we injected lines on (ITMX - ITMY) for MICH and PRM for PRCL. The optical gains were found to be

  • (ITMX - ITMY) to AS55_Q = 2.04e10 counts / m
  • PRM to REFL11_I = 1.66e12 counts / m

We added these into the C1:CAL_MICH_CINV and C1:CAL_PRCL_CINV filters on the calibration model.

PRMI-BHD handoff

After the lock was established and the alignment refined, we tried locking the homodyne phase angle using BH55_Q. We could not close the LO_PHASE loop with this sensor, so we resorted to using the BH44_Q. To do this we actually changed a few things:

  1. The BH44 whitening filter gain was lowered from +39 dB to +21 dB.
  2. Rotated the ND filter wheel to avoid saturation on the BH44 RFPD from 0.5 to 1.0. This corresponds to a change of ~33% in the incident BH44 light.

While we managed to lock BH44 by actuating on LO1, the lock was not very robust (typical lock acquistion using FM5 and FM8)... so we switched to actuating on AS4 which made the LO PHASE loop slightly more robust.

In preparation for handoff, we turned on a line at 211.11 Hz (MICH osc) and demodulated the BHDC_DIFF to estimate the sensing matrix element.

  • BHDC_DIFF to AS55_Q ~ 25

After using this number to rescale the MICH_B error point, we handed off from MICH_A (under AS55_Q control) to BHDC_DIFF. Our lock stretch covered the following gpstimes:

  • PRMI-BHD lock = 1368235235 to 1368235264

During the lock stretch above, we managed to take some noise spectra for the calibrated MICH and PRCL displacements. Attachment #1 shows the comparison between PRMI and PRMI-BHD configurations.


Discussion cool

  • Looking closely into Attachment #1 you'll find a *BH44_Q trace in green. This trace represents the uncalibrated LO phase error point. We measured the UGF to be < 20 Hz (C1:HPC-LO_PHASE_GAIN=1.2), so above 20 Hz, this trace is shown just as a guide for the noise shape. It would seem that the MICH displacement in PRMI-BHD readout is getting its shape from this noise and therefore might be limited by it.

Next steps

  • Calibrate LO phase at BH44 and see if it scales as predicted by our calculations.
  • Optimize BH44 control; why can't we lock this phase for longer?
  • Noise budget.
Attachment 1: prmibhd_Screenshot_2023-05-16_01-51-52.png
prmibhd_Screenshot_2023-05-16_01-51-52.png
  17591   Sat May 13 11:37:41 2023 ranaUpdateGeneralStripTool -> NDscope: PSL on wall

As a test, I have replaced the PSL StripTool on the north wall of the control room with a nearly equivalent NDScope display. It is plotting the real time second trend along with min/max values. Let's try this out for awhile, and if its bad we can just close the window.

FYI, the striptool machine, zita, has a mis-configured network setup so that its not getting a useful nameserver. So for now, its running on allegra through ssh -Y. We should put the zita network thing on some sort of issue tracker. Maybe we can have tickets and issues like various companies use for tech support? This could cover all of our main lab work and help us keep track of little problems like this.

Attachment 1: Screenshot_2023-05-13_11-45-47.png
Screenshot_2023-05-13_11-45-47.png
  17590   Thu May 11 12:05:24 2023 ranaUpdateBHDUpdated PRMI AS55+REFL11 noise budget

Is the A2L coming from the optical lever feedback? If so, we can make a 30 Hz ELP to cut it off by 60 Hz.

  17589   Wed May 10 17:11:34 2023 PacoUpdateASCFPMI differential ASC - I

I began commissioning the AS WFS for (PR)FPMI configurations

The main goal of this work is to close some ASC feedback loops during FPMI, or PRFPMI configurations to make locks more robust and their acquisition reliable. The running hypothesis is that angle to length coupling of the differential arms is increasing the rms and therefore straining the LSC controls... Today I began working on this, and unfortunately didn't make a lot of progress due to hardware hurdles, but the nominal plan is:

  • Stop IMC WFS and hijack the RFQPD heads for sensing -- WORK IN PROGRESS
    • We can temporarily break the IMC WFS loops to work on this upgrade, but in the future we would like to have at least one head (WFS1 or WFS2?) + transmission QPD to control IMC while the other head is our AS WFS.
    • Anchal already commissioned the AS WFS electronics and models. He also installed a couple of flipper mirrors at the Vertex table that pick the whole AS beam (otherwise going into the AS camera and ASDC, AS55, and AS110 PDs).
      • While this allows single ARM cavity ASC, we can't control MICH using our AS55 error signals.
      • I found that a beam was picked off along the AS110 path for another WFS head (currently unused), so after moving the head out of the way, I installed a mirror that sends it all the way to the IMC WFS heads area. The configuration before and after is summarized in Attachments #1-3.
      • The beam will need profiling again, but this doesn't destroy any of Anchal's ARM ASC alignment, just blocks it so we can recover it, so I think it's a good solution.
  • Lock FPMI and measure sensing matrix -- PENDING
  • Close loop, aim for gain of 10 at 10 Hz, then roll-off (e.g. IMC WFS) -- PENDING (RXA: G = 10 @ 1 Hz. G = 10 @ 10 Hz would be too spicy)

DESCRIPTION OF ATTACHMENTS

Attachment #1 before the changes were done

  • The red beam path indicates the nominal AS beam path
  • The purple beam path indicates the AS110 beam path.
  • The blue beam path indicates the unused, AS WFS (?) beam path.

Attachment #2-3 after the changes were done

  • The blue beam indicates the newly picked AS WFS beam path.
  • The yellow beam indicates the ARM AS WFS alternative path (installed by Anchal, controlled by flipper mirrors).
Attachment 1: PXL_20230510_233850237~2.jpg
PXL_20230510_233850237~2.jpg
Attachment 2: PXL_20230511_000946357~2.jpg
PXL_20230511_000946357~2.jpg
Attachment 3: PXL_20230511_000941172~2.jpg
PXL_20230511_000941172~2.jpg
  17588   Wed May 10 11:49:34 2023 YehonathanUpdateBHDUpdated PRMI AS55+REFL11 noise budget

I added input noises and angle to length coupling to the noise budget.

I added ADC and whitening filters noise contributions. The ADC noise is assumed to be 1uV/sqrtHz and the whitening noises were measured before in elog 17582. I use the measured whitening filter (elog   17584 ) to get the signal referred noise and calibrate.

The angle-to-length coupling is computed by taking the suppressed OpLev noise spectra of ITMX, ITMY, and BS and converting them to length noise by using the recently measured coupling coefficients in ELOG   17583 

Attachment 1: Quick_PRMI_noise_budget.pdf
Quick_PRMI_noise_budget.pdf
  17587   Tue May 9 21:02:55 2023 RadhikaUpdateALSXEND green PDH controller

XAUX laser locked with Moku:Go controller

The analog zeros and poles used to design this filter were:

zeros = [-18849.55592154, -18849.55592154]
poles = [-125.66370614, -238.76104167, -100530.96491487]
gain = 3000

Attachment 1 shows the resulting digital SOS filter (sampling rate: 3.9 MHz) compared to the measured uPDH servo transfer function (loop closed). The filter design was loaded on the Moku:Go.

Lock acquisition

I locked the AUX laser with the uPDH servo box and maximized its transmission to ~0.8. I then fed the Moku digital filter output to the PZT and the laser was able to catch lock. However, the max green transmission I could achieve using the Moku controller was 0.5. Attachment 2 is a screenshot of the green transmission ndscope during a lock sequence.

I measured the OLTF of the loop by injecting an excitation at the error point. An SR560 was used to sum the error signal with the excitation. The Moku multi-instrument mode was configured with the Frequency Response Analyzer and Digital Filter Box; it was able to source the excitation and take a transfer function measurement of error signal / (error signal + excitation), while keeping the loop closed.

The OLTF measurement [Attachment 3] points to a loop UGF of ~4 kHz, and phase margin of ~70 deg. An optimal controller would be able to boost the gain around the UGF without changing the phase too much (lag compensator)?

Attachment 1: PDHservoTF_eyeballZerosPoles.pdf
PDHservoTF_eyeballZerosPoles.pdf
Attachment 2: IMG_4721.JPG
IMG_4721.JPG
Attachment 3: XEND_AUX_Moku_OLTF.pdf
XEND_AUX_Moku_OLTF.pdf
  17586   Tue May 9 12:06:35 2023 RadhikaUpdateALSXEND green PDH controller

[Mayank, Radhika]

I retook a transfer function measurement of the uPDH servo closed-loop (using the SR560 to simulate a cavity pole) [Attachment 1]. While some coherence is lost at low frequencies, the servo does not appear to be saturating. Moving forward this measurement is used to design a digital filter that can replicate the uPDH servo box response. *Note: for now the chosen sampling frequency for the discrete filters is 61.04 kHz, the lowest sampling frequency setting of the Moku:Go.

We performed a low-order fit of the TF using vectfit. Vectfit always seems to return 1 more zero than pole - this results in an "improper" transfer function that causes any transformation to the z-domain to fail. Mayank took the fitted zeros and poles from vectfit and manually removed one of the zeros. After transforming the zeros and poles to the z-domain (using control.matlab.c2d), we noticed multiple resonances around 100 kHz that reached 10-20 dB. We decided to estimate poles and zeros by eye instead of using vectfit. 

2 zeros and 2 poles were selected by eye to get an estimated fit in the s-domain. Using continous-to-discrete transforms (tried scipy.signal.bilinear and control.matlab.c2d) resulted in unstable controller responses. Attachment 2 shows the original TF measurement with the designed analog filter and the resulting digital filter. The orange 'x's and 'o's mark the poles and zeros used. The digital filter contains many high-frequencies resonances, the most significant at the sampling frequency, 61.04 kHz, reaching 20 dB. Next we tried to manually load the analog ZPK coefficients into Foton. This resulted in the same digital filter as the python s-domain to z-domain functions [Attachment 3].

**UPDATE** Now looking back it's clear that the high-frequency response is limited by the sampling rate. I will redo this for the highest Moku:Go sampling rate of 3.9 MHz.

Attachment 1: PDHservoTF.pdf
PDHservoTF.pdf
Attachment 2: PDHservoTF_eyeballZerosPoles.pdf
PDHservoTF_eyeballZerosPoles.pdf
Attachment 3: eyeball_uPDH_fit.pdf
eyeball_uPDH_fit.pdf
  17585   Tue May 9 11:32:04 2023 YehonathanUpdateBHDWhitening TF measurements

We forgot to take a reference TF measurement by looping the SR785 on itself using the same BNC cables used for the actual measurement. I took this measurement today (attachment 1). As can be seen, there is a significant delay in the SR785 + cables themselves.

I also retook some measurements on the AS5_I whitening channel for various gains. Being careful with the excitation level and the channel range on the SR785 to avoid saturation I was also able to see low-frequency gains higher than 18dB so that problem is gone too. The results are shown in attachment 2 with the reference phase subtracted from the measurements.

Attachment 1: Reference_TF.pdf
Reference_TF.pdf
Attachment 2: Whitening_TFs_AS55_I.pdf
Whitening_TFs_AS55_I.pdf
  17584   Mon May 8 17:05:30 2023 YehonathanUpdateBHDWhitening TF measurements

{Mayank, Yehonathan}

We measured today the TFs of the whitening boards. We measured in particular REFL11 I/Q and AS55 I/Q channels using SR785.

There seems to be an issue with turning on whitening gain bigger than 18dB. In all our measurements, when the whitening filter was off the TF was flat and had the right gain. However, when we turned the whitening on, the measured TFs for gains higher than 18 dB would like exactly like as if the whitening gain was 18 dB. This happened in all channels that were measured and across two separate whitening filter boards.

Also, it was hard to measure both low and high-frequency parts of the TFs when the gain was high. The gain difference should be normally 40 dB but for higher gains it seems smaller. We verified that at higher gain level the high-frequency response was dependant on the ecxitation level meaning we had some saturation there.

 

Attachment 1: Whitening_TFs_REFL11_I.pdf
Whitening_TFs_REFL11_I.pdf
Attachment 2: Whitening_TFs_REFL11_Q.pdf
Whitening_TFs_REFL11_Q.pdf
Attachment 3: Whitening_TFs_AS55_I.pdf
Whitening_TFs_AS55_I.pdf
Attachment 4: Whitening_TFs_AS55_Q.pdf
Whitening_TFs_AS55_Q.pdf
  17583   Thu May 4 17:14:15 2023 PacoSummaryLSCPRMI (MICH and PRCL) calibrated displacement and BS angle to MICH couplings

[Mayank, Paco]

We calibrated PRCL displacement using MICH and estimated (?) BS angle to MICH couplings

Expt 1 - calibration of displacements for PRMI

We injected calibration lines at four different frequencies (211.11, 313.13, 575.17 and 1418.93) into (ITMY) under PRMI configuration and recorded the calibrated MICH displacement (C1:CAL-MICH_W_OUT) and uncalibrated PRCL displacement (C1:CAL-PRCL_W_OUT). The exctation amplitudes were (100,150,250 and 1000 cts) respectively.

The ratio of intensities of calibration lines peak in the two configurations are used to caliberate the PRCL displacement (C1:CAL-PRCL_W_OUT). Basically, if MICH = Lx - Ly and PRCL = Lp + 0.5(Lx + Ly), then PRCL displacement from Ly is half of MICH displacement.

The estimated optical gain (C_INV) was 6.41e-11 cts/m, which we added into FM3 of PRCL calibration filter bank. Attachment #1 shows the calibrated displacement spectra.

Expt 2 - angle to length couplings

After calibrating PRCL displacement, we tried to estimate the angle to length coupling from BS PIT/YAW to MICH under the PRMI configuartion. We started by injecting a broadband guassian noise (3000 cts amplitude in the 7-17 Hz band for PIT, and 17-27 Hz band for YAW) through an 8th order Cheby II filter with 80 dB attn to shape the noise before injecting into BS ASCPIT/ASCYAW.

We then looked at the increased displacement noise in the calibrated MICH and PRCL displacements. Assuming a naive 1 urad/count / f^2 of BS angular actuation strengths in both PIT/YAW, we estimate the couplings to be:

  • BS PIT to MICH = 0.12 nm / urad @ 13 Hz
  • BS YAW to MICH = 0.19 nm / urad @ 22 Hz
    • BS PIT to PRCL = 0.38 nm / urad @ 13 Hz
  • BS YAW to PRCL = 0.08 nm / urad @ 22 Hz

Seems a bit too large... blush bogus? Attachment #2 summarizes these measurements.

Attachment 1: calibrated_MICHPRCL_dispScreenshot_2023-05-04_18-46-05.png
calibrated_MICHPRCL_dispScreenshot_2023-05-04_18-46-05.png
Attachment 2: BSASC_noise_toMICHScreenshot_2023-05-04_18-59-32.png
BSASC_noise_toMICHScreenshot_2023-05-04_18-59-32.png
  17582   Wed May 3 18:40:50 2023 YehonathanUpdateBHDWhitening noises measurements

{Mayank, Yehonathan}

We measured the noise at the WF1 (REFL11) and WF2 (AS55) boards at the LSC rack with and without whitening filter. We switch the filter on and off by switching off and on the unwhitening in the PDs filter bank.

Attachment 1 shows the measurements.

Attachment 2 shows the ratio between the noise with and without whitening filter. I also plot the inverse of the unwhitening MEDM filter (all the unWhite filters were the same). I tune the gain of that filter to match the ratio of the AS55 whitening noises.

This is because I couldn't match the ratio of the REFL11 noises.

Moreover, the overall gain doesn't make sense to me. AS55 whitening has a gain of 24db and REFL11 has a gain of 18db. I'm not entirely sure where these values should show up. Also seems like REFL11 whitening has more gain than AS55 whitening. Will have to investigate more tomorrow.

Attachment 1: Whitening_noises.pdf
Whitening_noises.pdf
Attachment 2: Whitening_noises_on_off_ratios.pdf
Whitening_noises_on_off_ratios.pdf
  17581   Wed May 3 16:24:07 2023 PacoSummaryLSCAttenuated BHD RFPD paths

[Yuta, Paco]

We attenuated RFPD BHD paths to prevent saturation

To prevent saturating RPFDs, we added ND filters along BH44 and BH55 paths. We calculated the optimal filters to be used by taking into account the previously measured DC levels at both RFPDs as well as their DC transimpedances.

BH55 (S/N 117, former POP55) DC transimpedance is reported to be 10010 V/A. Assuming a responsivity of 0.8 A/V, we expect 8008 V/W at DC. When PRMI is locked, we have measured 10 mW of light incident on this PD such that it gets saturated. Then we installed an OD ~ 1 filter to drop the power to 1 mW such that at most we get 8 VDC (< 15 VDC rail).

BH44 (S/N 115, former unknown) DC transimpedance is estimated to be 645 V/A. Assuming 0.8 A/V, we expect 516 V/W at DC. This RFPD gets 27 mW when PRMI is locked, giving a slightly saturated PD> Then we installed an OD ~ 0.5 filter to drop the power to ~ 8 mW such that we at most get 4 VDC (< 15 VDC rail).

  17580   Wed May 3 15:06:44 2023 yutaSummaryLSCPOP attenuation and PRMI PRG estimate

[Paco, Yuta]

Measured power recycling gain at POP is 10(2), consistent with our expectation.

We measured power at POP with ITMY single bounce and estimated power recycling gain in PRMI.
As POP RFPD (Thorlabs PDA10CF; used for POPDC, POP22, POP110) was saturating, we attenuated the input power by OD2.5 ND filter.
Power recycling gain was estimated to be 10(2), roughly consistent with our expectation of 13.2 (40m/17532).

What we did:
 - Realigned POP path in ITMX table after aligning the IFO. It turned out that when POP power measurements were done in 40m/17532, POP was not well aligned.
 - We measured the power with ITMY single bounce at POP right after the viewport and in front of POP RFPD.
 - We also measured counts in C1:LSC-POPDC_OUT under ITMY single bounce, PRMI carrier locked, and MICH locked with PRM misaligned, with different ND filters.

Results:

IFO configuration Where Measured power [mW] C1:LSC-POPDC_OUT [counts]
ITMY single bounce POP total 0.224(5) Expected 0.240 [a] N/A
  POP RFPD (no ND filter) 0.108(3) 437(2)
  POP RFPD (OD1) N/A 29.9(1), which is 7.4 uW [b]
  POP RFPD (OD2+OD0.5) N/A 2.4(5), which is 0.6 uW [b]
PRMI carrier POP RFPD (no ND filter) N/A 13160 (saturated)
  POP RFPD (OD1) N/A 12300(300) (saturated)
  POP RFPD (OD2+OD0.5) N/A 1600(300), which is 0.40 mW [b]
MICH POP RFPD (OD2+OD0.5) N/A 9.0(6), which is 2.2 uW [b]

 - Estimated power recycling gain is 1600(300) / (9.0(6) / 5.637%) = 10(2) .

Expected values:
 - Expected power using PSL output of 890 mW (measured in elog 40m/17390) under ITMY single bounce at POP is the following, and is consistent with the measurement.
[a] 890 mW * 0.9 (IMC transmission?) * 5.637%(PRM) * (1-2.2%)(PR2) * 50%(BS) * (1-1.384%)(ITMY) * 50%(BS) * 2.2%(PR2) = 0.240 mW

 - Calibration for C1:LSC-POPDC_OUT into power at POP RFPD is
[b] 437 counts / 0.108 mW = 4.0e3 counts/W.

 - Thorlabs PDA10CF has a transimpedance gain of 1e4 V/A and output range of 0-5 V. So, the saturation happens at 5 V / (1e4 V/A * 0.8 A/W) = 0.625 mW. We ended up attenuating POP RFPD with OD2.5 to make it not saturating (0.4 mW on the PD with PRMI carrier lock).

  17579   Wed May 3 12:11:52 2023 YehonathanUpdateBHDUpdated noise budget with measured noise and OLTF

{Paco, Yehonathan, Yuta}

Paco and Yuta locked PRMI carrier and I took the MICH OLTF measurement (attachment 1).

I downloaded 300secs of C1:LSC-MICH_IN1_DQ from when the PRMI was locked yesterday and calibrated it with the OLTF. I plot it together with the noise budget (attachment 2).

Attachment 1: PRMI_carrier_MICH_OLTF.pdf
PRMI_carrier_MICH_OLTF.pdf
Attachment 2: Quick_PRMI_noise_budget.pdf
Quick_PRMI_noise_budget.pdf
  17578   Tue May 2 17:33:53 2023 MayankSummaryLSCLocked PRMI in carrier for an hour with LO phase controlled using BH55

[Mayank, Radhika, Paco]

We locked PRMI for a solid hour devil and controlled LO phase angle using BH55_Q at higher power.

After Radhika aligned the IFO for us, and recovered the PRMI flashing (using REFLDC), we attempted a PRMI lock. After a few trials we succeeded.

Control parameters: see Attachment #1, basically REFL11_I to PRCL, and AS55_Q to MICH (error points) and actuation as previous locks with PRCL to PRMand MICH to 0.5 * BS - 0.33 * PRM.

The gains are slightly different, and in particular PRCL gain was increased from -0.07 to -0.09 after an OLTF estimated the UGF could be increased to > 120 Hz (Attachment #2 shows the measured OLTF) Do note we ended up disabling the FM1 on PRCL LSC filter bank (a boost) because we thought the loop was unstable when it got triggered ON. Finally, we took a quick noise spectrum of PRMI, and we have yet to calibrate it.

We also managed to reduce the AS_DC level from 0.4 to 0.1. We first tried to add an offset to MICH error point but the trick was to align the ITMX ITMY differential yaw.

Lock start Time: 1367107965 --> 1367111565


While PRMI was locked, we quickly locked homodyne angle using BH55_Q. For this the demod angle was optimized from -60 deg to 55.374 deg. The lock was acquired using FM5 and FM8 with a gain of -0.75. Attachment #3 shows the "calibrated" noise budget of the LO phase under this configuration. The main difference with respect to the previous  budget is in the "RIN" which we now realize is not relative... therefore the increase in this budget. We will revisit this calibration later.


- Next steps

  • Re-calibrate LO phase noise with high power
  • Investigate BH44 control
  • Calibrate PRMI noise for budget
  • Estimate LO phase sensitivities at MICH vs PRMI

 

Attachment 1: Screenshot_2023-05-02_17-40-06.png
Screenshot_2023-05-02_17-40-06.png
Attachment 2: PRCL_OLTF_Screenshot_2023-05-02_18-10-28.png
PRCL_OLTF_Screenshot_2023-05-02_18-10-28.png
Attachment 3: PRMI_LO_phase_BH55_Q_Screenshot_2023-05-02_18-06-59.png
PRMI_LO_phase_BH55_Q_Screenshot_2023-05-02_18-06-59.png
  17577   Tue May 2 10:39:49 2023 JCConfigurationIMCBad Alignment

It took a while, but I was finally able to align IMC. It seems like WFS has been getting really whacky lately when we arent in the lab watching it angry. The picture attached has an arrow of where the beam spot was at this morning

Attachment 1: BB26DF83-5B6A-4BC1-A2A0-6E18E1A2B91E.jpeg
BB26DF83-5B6A-4BC1-A2A0-6E18E1A2B91E.jpeg
  17575   Mon May 1 16:51:20 2023 PacoSummaryASCIFO alignment in bad shape

[Yuta, Mayank, JC, Paco]

We fixed the IFO nominal alignment.


  1. Yuta and Mayank had worked all morning and into the early afternoon to try and recover alignment. We noticed a few things seemed off:
    1. ASS loops still wouldn't work.
    2. ITMX oplev loops were weirdly unstable, but we suspected this had to do with the recent HeNe replacement saga.
      1. Paco revisited the setup, suspecting a lens near the QPD was wreaking havoc. This was not the case, but the setup now resembles the previous one. Furthermore, an iris was placed in the input ITMX OpLev path.
    3. Using AUX (green) beams as a reference didn't really work well.
  2. We decided to check the SUS rack electronics where some noise measurements were carried out last week.
  3. We found out that between the DW boards from ITMX and ITMY to the coil driver units for ITMX and ITMY there were two crossed wires at least (Ch4).
  4. Yuta and Paco reconnected all channels between the ITMX DW to ITMX Coil driver and ITMY DW to ITMY Coil driver units,
    1. All wires seemed to be crossed when going up the rack... suspicious...
  5. The first damping test failed, and we realized placing an offset in ITMY coils affected ITMX, so the DW board units were probably flipped because they were probably mislabeled!
  6. Yuta and JC swapped the cables again, and we ran a coil by coil test before damping. After this ITMs were successfully damped, so the labels were corrected in the AI + DW boards to prevent this confusion in the future.

  • IFO alignment was recovered by Yuta and Paco.
    • ASS now works again!
  • JC and Mayank aligned OpLevs
    • Loops were closed and remained stable in ITMs.

Attachment #1 shows the alignment state at the end of this work.

Attachment 1: alignmentScreenshot_2023-05-01_17-06-41.png
alignmentScreenshot_2023-05-01_17-06-41.png
  17574   Mon May 1 14:45:48 2023 MayankSummaryLSCAttenuated BHD DC Beam

[Yuta, Mayank]

UPDATE: It turned out that the pair of 0.3 OD ND filters we used were not matched. So we replaced them with new 0.5 OD NENIR05A-C from thorlabs. Now both the photodiodes give similar count.

Counts Before OD After OD
C1:HPC-BHDC_A_OUT 114.5 counts 36.4 counts
C1:HPC-BHDC_B_OUT 111.5 counts 35.1 counts

 

 

The DC power incident on the PDs is 74 mW which may cause saturation. We attenuated the beam going to BHD_DC Photodiodes using ND filter of OD 0.3 which gives attenuation of 0.5. 

Attachment 1: 20230501_144152.jpg
20230501_144152.jpg
  17573   Mon May 1 08:57:45 2023 JCConfigurationIMCBad Alignment

I had to realign the IMC today. When I came in, it was very bad, not much flashing at all, I had to do it from scratch. CH01 Camera on MON7 in the control room is completely white. Did the camera go out over the weekend? I will come back to poke around later, I have headed over to WB for a moment and will be back soon. sitemapsi

Attachment 1: Screenshot_2023-05-01_15-55-33.png
Screenshot_2023-05-01_15-55-33.png
  17572   Fri Apr 28 20:56:06 2023 KojiSummaryASCIFO alignment in bad shape

I suppose ASS Y arm is using PR2/3 to align the beam to the arm.
Can't we have ASS PRM bring the beams to the center (or some defined places) of the PRM and PR2 by moving the TTs?

 

  17571   Fri Apr 28 20:17:37 2023 yutaSummaryASCIFO alignment in bad shape

[Mayank, Paco, Yuta]

IFO alignment is not good.
It seems like the input pointing drifted a lot during PRMI and noise measurements, and beam spot on both ITMY and ITMX are not good.
They are so off from the center (by about a beam size mainly in yaw) that ASS cannot handle.
Current situation is as attached (compare with good alignment in March 23 40m/17521).
Yarm ASS is not working, Xarm ASS is not working, POP is clipped, AS is clipped crying

Message: Always check the alignment from TTs using BHDC_A/B, and always check the arm alignment, even if you are only doing PRMI. (Follow the steps in 40m/17277)

Attachment 1: Screenshot_2023-04-28_20-16-19_Terrible.png
Screenshot_2023-04-28_20-16-19_Terrible.png
  17570   Fri Apr 28 18:40:49 2023 YehonathanUpdateBHDUpdated noise budget with some input electronic noises

{Mayank, Yehonathan}

Yesterday, we measured AS55 and REFL11 dark noises at the IQ demod boards outputs (attachment 1) using SR560+SR785 setup.

We also measured the whitening board noise of REFL11 using an improvised adapter (picture will come later). The measurement result is shown in attachment 2. Didn't have time to measure the whitening noise for AS55

Also, after realizing the Finesse model doesn't account for the REFL port attenuation I measure how much DC power at the REFL11 PD to be 0.8mW by aligning PRM and misaligning the rest of the optics.

For some reason, the power before the attenuation is only ~ 360mW. The Finesse model predicts around 700mW. Where is the rest of the light going?

I added the PDs Dark noises (using the recently measured IQ demod gains) and shot noise to the PRMI carrier noise budget (attachment 3). ADC and whitening noises coming soon.

Measurements and PRMI noise budget notebooks were uploaded to the 40m git.

Attachment 1: PDIQ_Demod_noises.pdf
PDIQ_Demod_noises.pdf
Attachment 2: Whitening_noises.pdf
Whitening_noises.pdf
Attachment 3: Quick_PRMI_noise_budget.pdf
Quick_PRMI_noise_budget.pdf
  17568   Wed Apr 26 18:41:56 2023 PacoUpdateOPLEV TablesITMX oplev laser died

[Paco, Koji, Mayank]

We replaced ITMX Oplev HeNe laser after last one died.

  • We noticed the ITMX HeNe laser died, so we couldn't have Oplev control. With Koji's help, we located a laser tube which had no labels and a sticker up front so we assumed this is new. We removed the faulty tube from the ITMX table and tested it using the display control unit (labeled 1). The laser head was 1103, and weirdly it turned on. We also located some control units in the Oplev cabinet which had been labeled "working with Display" from 2022.
  • After testing a few combinations we realized CARE NEEDS TO BE TAKEN TO NOTICE THE INTERLOCK SWITCH BEFORE TESTING ANY PAIR> it turned out our old controller was ok, but the interlock switch had fallen. We also noted the two spare units labeled "working with Display" were indeed operating correctly they just needed their interlock switches. Surprisingly the "new" laser head we planned to use as a spare is DEAD.
  • After assuming the old head was working and its controller just needed the interlock, we reinstalled it on the table. Unfortunately, after realigning the OPLEV beam we saw the SUM drop from 1500 counts (usually 15000) to 0. So the laser head was indeed faulty! + WE HAVE NO MORE 1103 SPARE HEADS!
  • We couldn't find another 1103 spare head, so we opted for the 1125 along with its corresponding controller. This worked fine, except the unit is thicker and the whole beam path changed by ~ 1 cm in height. Mayank and I restored the alignment and centered the ITMX oplev. A photo of the changes made to reach the final setup are in Attachments #1-2. Most notably we had to lower the big pedestals where the HeNe laser head is resting to accommodate this height difference.
  • Finally the laser beam diameter is different, so we removed the "collimating" lens immediately after the HeNe head. This had the effect of restoring the beam path, but we placed a shorter focal length lens between the last steering mirror and the QPD> this means the previous ITMX oplev calibration is now bogus.
Attachment 1: PXL_20230427_004834747.jpg
PXL_20230427_004834747.jpg
Attachment 2: PXL_20230427_013642978.jpg
PXL_20230427_013642978.jpg
  17567   Wed Apr 26 12:59:42 2023 YehonathanUpdateBHDUpdated noise budget with output electronics

I included the output electronic noises into the PRMI carrier noise budget (attachment 1).

The coil driver noise was calculated using the Johnson noises of the coil driver resistor:

PRM 430 ohm

BS 100 ohm

ITMX/Y 400 ohm

For the dewhitening noises I use the measurements from yesterday. As expected fro yesterday's measurements, the ITMX dewhitening noise is dominating. For the coil driver gain I use the recently measured actuation calibration (elog   17522 ) to extract it. I find that these gain values:

PRM 1.009

BS 1.333

ITMX/Y 0.24

For the DAC noise I assume 1uV/sqrtHz and use the simDW filters from the coil outputs MEDM screens as the DW filters TFs.

Next:

1. Break down input noises.

2. Measure how much light is reaching REFL11 to correct the sensing matrix and get the right shot noise.

Attachment 1: Quickl_PRMI_noise_budget.pdf
Quickl_PRMI_noise_budget.pdf
  17566   Wed Apr 26 12:05:10 2023 RadhikaUpdateALSXEND green PDH controller

Tl;dr: Tried to replace of XEND green PDH servo controller with Moku template IIR filter, designed to match PDH servo frequency response. The green laser did not catch lock with this filter.

Attachment 1 plots the measured TF of the PDH servo controller, with boost on and the gain knob set to 7.22 (the current lock configurations). It also plots an 8th order Chebyshev type II low-pass filter, with cutoff frequency and scale chosen to best match the data. (8 was the highest order filter that could be represented by 4 second-order-sections, the maximum allowed by the Moku.) I wanted to test if the XAUX PDH lock could be maintained using this filter as the controller.

The phase of the Chebyshev II filter does not seem to be a good fit to the data, but I wanted to see how far we could get using a template filter already designed for discrete time, and with a magnitude frequency response that approximates the servo. This would bypass having to perform a bilinear transform from the s-domain to the z-domain, which can raise more complications.

The PDH error signal (mixer output) was split and sent to the Moku (input 1) and to the PDH servo input. Closing the loop with the Moku filter output, the green laser was not able to catch lock. Attachment 2 shows the Moku:Go Digital Filter Box configurations, as well as the traces comparing output of the filter and the output of the PDH servo. The red trace is the output of Moku filter, and the blue trace is the output of the PDH servo (input 2) with the loop open (nothing feeding back to laser PZT). The input gain of the filter module was chosen to match the amplitudes of the two control signals. Qualitatively, the filter output contains higher frequency components and preserves the odd polarity of the PDH error signal, compared to the servo output. 

I then tried to directly fit the PDH servo TF data. I fit the (analog) poles and zeros of the TF using vectfit. In theory, using a bilinear transform can convert the analog zpk TF to digital zpk, with some frequency pre-warping required. However, vectfit did not return a "normal" transfer function, defined as having at least as many poles as zeros. This caused the bilinear transform to fail.

Next, I will need to use a different fitting package (perhaps IIRrational) to obtain a nicer TF fit, in normal form. Then I can attemp the bilinear transform, confirm it preserves the desired frequency response, and test it out with the Moku:Go.

Attachment 1: PDHservoTF_chebyIIfilter.pdf
PDHservoTF_chebyIIfilter.pdf
Attachment 2: Screenshot_2023-04-25_at_09.28.05.png
Screenshot_2023-04-25_at_09.28.05.png
  17565   Wed Apr 26 11:27:49 2023 PacoSummaryBHDLO/MI(DARM) signal strength comparison between the configurations with finesse

I'm checking Koji + Yuta's not-so-naive calculations using finesse.

  Michelson PRMI carrier PRMI sideband PRFPMI
max(BH44) [W/m] 0.61 @ 90 deg 235.76 @ 90 deg    
max(BH55) [W/m] 4.55 @ 0 deg 1539.67 @ 0 deg    
max(BHD_DIFF) [W/m] 35550 10656140    

PRMI Carrier resonant:
- The HD phase signal @44 MHz is estimated to be 386.5 times stronger.
- The MICH signal @BHD_DIFF is estimated to be enhanced by a factor of 299.75.

PRMI Sideband resonant:
- The HD phase signal @44MHz is estimated to be () stronger.
- The MICH signal @BHD_DIFF is estimated to be suppressed by a factor of

PRFPMI: Now the carrier and the 11 and 55MHz sidebands are resonant.
- The HD phase signal @44MHz is estimated to be ().
- The MICH signal @BHD_DIFF is estimated to be the same level as the one with PRMI Carrier resonant.

  17564   Wed Apr 26 09:37:10 2023 PacoUpdateBHDIQ demod board gains for REFL11 and AS55

We measured the IQ demodulation board gains for REFL11 and AS55.

To do this, we replaced the PD input on the demod board with an RF signal at near the nominal frequencies of 11.066195 MHz and 55.330975 MHz using a Marconi 2024A identical to the one which sources the PM sidebands in our PSL. Even though we matched the modulation frequencies we found the two marconis were in practice offset by ~ 3 Hz. After tuning the frequency around a bit, we managed to get them to within 450 mHz.


REFL11

We started with REFL11 IQ demod board. After sourcing 11.066198 MHz into the PD input port, we took the I and Q outputs and looked at them using an osciloscope. We measured the Vpp levels on both as well as the Marconi output. The resulting levels were

  • Source = 44.4 mVpp, I = 8.8 mVpp and Q = 10.8 mVpp ==> Gains are therefore 0.19 and 0.24. The amplitude gain of this board is sqrt(0.19 ** 2 + 0.24 ** 2) = 0.153. This is in stark disagreement with the wiki. Has the wiki finally failed us?

AS55

We then moved on to AS55 IQ demod board. After sourcing 55.330975 MHz* into the PD input port, we took the I and Q outputs and looked at them using an osciloscope. We measured the Vpp levels on both as well as the Marconi output. The resulting levels were

  • Source = 16.8 mVpp, I = 50.8 mVpp and Q = 56.3 mVpp ==> Gains are therefore estimated to be 3.3 and 3.7. The amplitude gain of this board is sqrt(3.3**2 + 3.7**2) = 4.74. This is in slight contrast with the previously measured gain of 2.8, but we think a factor of 2 may have been misplaced in either calculation since one typically estimates AMP = 2 * sqrt(I**2 + Q**2).

* Note that in the second test, we didn't match up the frequency, which caused I and Q outputs to have significant gains (instead of just I).

  17563   Tue Apr 25 21:21:03 2023 YehonathanUpdateBHDDewhitening noises

{Mayank, Paco, Yehonathan}

Dewhitening noise curves were taken using SR785+SR560 for the PRMI noise budget. One representative channel was measured at each board, suspensions were tripped before work was done. The input pins to the dewhitening boards were shorted using an exposed ribbon cable.

At each board, the measurement was taken with and without dewhitening filter on. The toggling of the dewhitening filter was done by turning on and off the SimDW filters at the coil filter screen of each suspension.

Attachment 1 summarizes the results.

ITMX dewhitening noise is much higher than the rest.

ITMY measurement turned out to be bogus since we mostly measured dark noise. The reason we made the gain so low in that measurement is that it was saturating the SR560 whenever we used gain>1.

Attachment 1: De-whitening_noises.pdf
De-whitening_noises.pdf
  17562   Tue Apr 25 17:06:17 2023 AnchalUpdateALSDFD demod normalized by amplitude

I modified the analysis to correct for any affects due to Anti-Aliasing or Anti-Imaging filters, and I also found a insignificant error on how I was undoing the suppression due to MICH loop in the MICH data. I also propagated the calibration in MICH method better. Attached are the updated results. The upward swing is still present.

Also, last night, Koji and I looked into any frequency dependent deviation in sensing arm length between POY11 and BEATY_PHASE (using DFD+Phase tracker) This was done by locked the YARM to the main laser and locking YAUX to the YARM, sending excitationa at C1:SUS-ETMY_POSCAL_EXC and taking transfer function between C1:LSC-YARM_IN1 and C1:ALS-BEAT_Y_FINE_PHASE_OUT. This transfer function was flat upto about 600 Hz and the deviation from there to 2000 Hz was expected based on limited bandwidth of the phase tracker. I don't have the plot to attach, someone should redo this quick measurement to save the data.
Interestingly, the same measurement when done with  C1:LSC-DARM_IN1 in FPMI configuration did not show a flat response. This is can mean that the DARM strain relationship with the beatnote frequency deviation is not a simple constant factor and/or depends on DARM or CARM OLTFs. I leave my remarks on this project here for the baton to be picked up by others in future. I unfortunately only have this much time to contribute to FPMI calibration.

Attachment 1: ITMYActCal.pdf
ITMYActCal.pdf
Attachment 2: ITMYActCalTS.pdf
ITMYActCalTS.pdf
  17561   Tue Apr 25 15:51:21 2023 JcUpdateIMCIMC has been tripping

It has happened multiple times today that IMC has tripped on its own. Yehonathan and I have had to come back to manually lock IMC multiple times.


Wed Apr 26 10:24:07 2023 [EDIT]

[Paco] I aligned the MC by hand, let it run locked for 30 minutes without angular controls, and then switched on the WFS loops yesterday at ~ 6 PM. IMC has been locked ever since.

Attachment 1: Screenshot_2023-04-25_15-55-33.png
Screenshot_2023-04-25_15-55-33.png
  17560   Mon Apr 24 19:11:20 2023 KojiSummaryBHDLO/MI(DARM) signal strength comparison between the configurations

Yuta and I had a discussion last week about the signal strength between the configurations. Here are some naive calculations.
=== Please check the result with a more precise simulation ===


Michelson: Homodyne (HD) phase signal @44MHz is obtained from the combination of LO11xAS55 and LO CAxAS44. SBs at AS rely on the Schnupp asymmetry, the signal is weaker than the one with a single bounce beam from an ITM.

PRMI Carrier resonant:
- Despite the non-resonant condition of the sidebands, the HD phase signal @44MHz is expected to be significantly stronger (~x300) compared with the MI due to the resonance of the carrier and the 44MHz sidebands (the 2nd-order SBs of 11 and 55) in the PRC. Thus, the LO CAxAS44 term dominates the signal.
- The MICH signal @55MHz is enhanced by the resonant carrier by a factor of ~5.5, in spite of the non-resonant 55MHz SBs.
- The MICH signal @BHD is enhanced by the resonant carrier by a factor of ~300. This is the comparable phase sensitivity to PRFPMI case.

PRMI Sideband resonant:
- Despite the non-resonant condition of the carrier, the HD phase signal @44MHz is expected to be even stronger (~x400) compared with the MI due to the resonance of the 11MHz and 55MHz sidebands in the PRC. Thus, the LO11xAS55 term dominates the signal.
- The level of the MICH signal @55MHz is expected to be comparable to the one with PRMI carrier resonant as the resonant condition for the CA and 55MHz SBs are interchanged.
- The MICH signal @BHD is expected to be negligibly small due to non-resonance of the carrier.

PRFPMI: Now the carrier and the 11 and 55MHz sidebands are resonant.
- The HD phase signal @44MHz is expected to be the same level as the SB resonant PRMI, and the LO11xAS55 term dominates the signal.
- The level of the MICH sensitivity @AS 55MHz shows x300 of the MICH signal of the MI and x50 of the MICH with PRMI.
- The MICH signal @BHD is going to be the same level as the one with PRMI Carrier resonant.
-
The DARM signal shows up at the dark port signal enhanced by x300 from the MICH level due to the finesse of the arms.



Simple assumptions
1) PRM has a transmission of TPRM = 0.05
2) PRG is limited by the transmission of PR2 (TPR2=0.02 per bounce).
    If the IFO is lossless, PRG is 25 (i.e. theoretical maximum). In reality, the IFO loss is 2~3% -> PRG is ~15.
    The asymmetry of 30mm has a negligible effect.
3) For the anti-resonant fields, APRG is ~TPRM/4 = 0.0125
4) Arm finesse is 450. Therefore the phase enhancement factor N is ~300.
5) Modulation depth is ~0.1. J0=1, J1=0.05, J2=0.00125
6) Sideband leakage by the asymmetry is ɑ=l_asym wm / c = 0.008 for 11MHz and 5ɑ for 55MHz.


Single Bounce

The numbers are power transmission to each port
   Carrier              11MHz                    55MHz
LO
TPRM TPR2 = 1.0e-3   J1^2 TPRM TPR2 = 2.5e-6  J1^2 TPRM TPR2 = 2.5e-6
AS TPRM/4    = 1.3e-2   J1^2 TPRM/4    = 3.1e-5  J1^2 TPRM/4    = 3.1e-5

LO phase @44MHz: LO 11 x AS 55 = Sqrt(2.5e-6 * 3.1e-5) = 8.8e-6


Michelson

   Carrier               11MHz                     55MHz                      44MHz
LO
TPRM TPR2 = 1.0e-3    J1^2 TPRM TPR2 = 2.5e-6   J1^2 TPRM TPR2   = 2.5e-6
AS TPRM ε^2  = 0.05 ε^2  ɑ^2 J1^2 TPRM  = 8.0e-9   25 ɑ^2 J1^2 TPRM = 2.0e-7  16 ɑ^2 J1^4 TPRM = 3.2e-10

LO phase @44MHz: LO 11 x AS 55 = Sqrt(2.5e-6 * 2.0e-7)  = 7.1e-7
                 LO CA x AS 44 = Sqrt(1.0e-3 * 3.2e-10) = 5.7e-7

AS MICH  @55MHz: AS CA x AS 55 = Sqrt(0.05 * 2.0e-7) ε  = 1.0e-4 ε
AS MICH  @BHD:  LO CA x AS CA = Sqrt(1.0e-3 * 0.05) ε  = 7.1e-3 ε


PRMI (Carrier Resonant)

   Carrier              11MHz                     55MHz                      44MHz
LO PRG TPR2 = 0.3       J1^2 APRG TPR2 = 2.5e-7   J1^2 APRG TPR2   = 2.5e-7  J1^4 PRG TPR2 = 1.9e-6
AS PRG ε^2  = 15 ε^2    ɑ^2 J1^2 APRG  = 8.0e-10  25 ɑ^2 J1^2 APRG = 2.0e-8  16 ɑ^2 J1^4 PRG = 9.6e-8


LO phase @44MHz: LO 11 x AS 55 = Sqrt(2.5e-7 * 2.0e-8)  = 7.1e-8
                 LO CA x AS 44 = Sqrt(0.3 * 9.6e-8)     = 1.7e-4
AS MICH  @55MHz: AS CA x AS 55 = Sqrt(15 * 2.0e-8) ε    = 5.5e-4 ε

AS MICH  @BHD:  LO CA x AS CA = Sqrt(0.3 * 15) ε       = 2.1 ε
 


PRMI (Sideband Resonant)

   Carrier               11MHz                     55MHz                      44MHz
LO APRG TPR2 = 1e-4      J1^2 PRG TPR2 = 7.5e-4    J1^2 PRG TPR2   = 7.5e-4  J1^4 APRG TPR2 = 6.3e-10
AS APRG ε^2  = 5e-3 ε^2  ɑ^2 J1^2 PRG  = 2.4e-6    25 ɑ^2 J1^2 PRG = 6.0e-5  16 ɑ^2 J1^4 APRG = 3.2e-11


LO phase @44MHz: LO 11 x AS 55 = Sqrt(7.5e-4 * 6.0e-5)  = 2.1e-4
                 LO CA x AS 44 = Sqrt(1e-4 * 3.2e-11)   = 5.7e-8
AS MICH  @55MHz: AS CA x AS 55 = Sqrt(5e-3 * 6.0e-5) ε  = 5.5e-4 ε

AS MICH  @BHD:  LO CA x AS CA = Sqrt(1e-4 * 5e-3) ε    = 7.1e-4 ε


PRFPMI

   Carrier             11MHz                     55MHz                      44MHz
LO PRG TPR2 = 0.3      J1^2 PRG TPR2 = 7.5e-4    J1^2 PRG TPR2   = 7.5e-4   J1^4 APRG TPR2 = 6.3e-10
AS PRG ε^2  = 15 ε^2   ɑ^2 J1^2 PRG  = 2.4e-6    25 ɑ^2 J1^2 PRG = 6.0e-5   16 ɑ^2 J1^4 APRG = 3.2e-11


LO phase @44MHz: LO 11 x AS 55 = Sqrt(7.5e-4 * 6.0e-5)  = 2.1e-4
                 LO CA x AS 44 = Sqrt(0.3 * 3.2e-11)    = 3.1e-6
AS MICH  @55MHz: AS CA x AS 55 = Sqrt(15 * 6.0e-5) ε    = 3.0e-2
ε  ==> DARM@55MHz 9.0 ε
AS MICH  @BHD:  LO CA x AS CA = Sqrt(0.3 * 15) ε       = 2.1 ε     ==> DARM@BHD   6.3e2 ε


  17559   Mon Apr 24 18:33:22 2023 AlexUpdateIMCBeam offset movement for MC1,2,3 in PIT and YAW from dither results

Mayank and I worked on finalizing the plots for the beam offset from the dithering test done in 17552. Plotted in attachment 1 are the beamspot demodulated signals from MC_F_DQ which are averaged over 1 second each (blue) for YAW and PIT in MC1,2,3. The yellow line over each plot shows the 3 Hz lowpassed signal of the beamspot movement.

Additionally, we have seen no direct correlation to the WFS1 or 2 sensors due to the MC movements. This may be because the WFS display a complete signal that includes all changes in the cavity length due to the shaking of the mirrors. Thus, the signal (shown in red) of the WFS sensors will show a combined average of movement from all 3 dither lines.

Attachment 1: beam_spot_time_series.png
beam_spot_time_series.png
ELOG V3.1.3-